Consider the real vector space
A = {(a, b, c, d) I a, b, c, d belongs to R, 2a + 3b = c + d}.
Find dim (A). Also find two distinct subspaces
B1. and B2 of R⁴ such that
A direct sum B1= R⁴=A direct sum B2
1
Expert's answer
2020-10-14T14:16:29-0400

The given vector space is

"A=\{(a,b,c,d ) : a,b,c,d \in \R ,2a+3b=c+d\}"

"=\{ (a,b,c,2a+3b-c):a,b,c \in\R\}...........(1)"

"=\{ (a,b,2a+3b-d,d):a,b,d\in \R\}..............(2)"

Since , "A" is depending on three real number . Therefore "d(A)=3"

Consider the subset "B_1 \ and \ B_2 \ of \ \R^4" defined as follow,

"B_1=\{(0,0,0,d):d\in \R \}"

"B_2=\{ (0,0,c,0):c\in \R \}"

Clearly, "B_1" and "B_2" are subspace of "\R^4"

If "A" in the form (1) then Our claim is

"\R^4=A\bigoplus B_1"

Let "(a,b,c,d)\in \R^4" be any arbitrary element.

Then "(a,b,c,d)=(a,b,c,2a+3b-c)+(0,0,0,d-2a-3b+c)"

Clearly , "(a,b,c,2a+3b-c) \in A"

and "(0,0,0,d-2a-3b+c)\in B_1"

Hence , "\R^4=A+ B_1"

Now , My next claim is

"A\cap B_1=\{0\}"

Let "(x,y,z,w)\in A\cap B_1" be any arbitrary element then ,

"(x,y,z ,w)\in A" and "(x,y,z,w) \in B_1"

Therefore, "2x+3y=z+w \ and \ x=y=z=0"

"\implies w=0"

"\implies(x,y,z,w)=(0,0,0,0)"

Therefore , "A\cap B_1 =\{0\}"

Hence , "A\bigoplus B_1=\R^4"

Similarly , If "A" is written in the form (2) then "A\bigoplus B_2=\R^4"